OpenStudy (anonymous):

What equation results from completing the square and then factoring x^2+4x=5? A. (x+2)^2=7 B. (x+2)^2=3 C. (x+2)^2=1 D. (x+2)^2=9

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

what is half of 4?

OpenStudy (anonymous):

2...

OpenStudy (anonymous):

k right, and what is \(2^2\)?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ok then your answer is \[(x+2)^2=5+4\] or \[(x+2)^2=9\]

OpenStudy (anonymous):

Oh. Well that was easy. Thanks!

OpenStudy (anonymous):

lol yeah when you know how it is real real (real) easy